An equation of the form ypx)ygx)y, a 0,1 is called the Bernoulli equation If we divide by ya we get g(x) . у Чу + yay p(x
Answer #1

Given that a gltzay - y =0 @ Divide it by n we get sta ja ne go - compare it by yt play ý =3(4) oyo. :þcu = 3:30) - Rs. LetSdk IF e - = mina of en ② is . The genera solation 4. (8-F) = false) dr. a C - ] = 3 2 2 2 and the =-2 fede astu - s + c M५= 3 +c: Have u = sz yaxy = Suste multiply with a, wagt which is the general solution of giren ☺ consider y=&y-ayo compare w-sende wiesze eextc BEX, we we - 2EX Et ceex then u:52. Feste ēzex which is the generd solation of en